The width of a rectangle is one third the length. The perimeter is 80 yards. Find the width of the rectangle
Answer by Cromlix(4381)

Length = x
Width = x/3
Perimeter = 2*length + 2*width
80 = 2(x) + 2(x/3)
80 = 2x + 2x/3
Common denominator = 3
240 = 6x + 2x
8x = 240
x = 30
Therefore the width - x/3 = 30/3 = 10 units
Hope this helps.
:-)